OpenStudy (anonymous):

Two positive numbers have a difference of 8 and a product of 33. What are these numbers? 4 and 12 3 and 12 4 and 11 3 and 11

OpenStudy (nurali):

12 - 4 = 8 12 * 4 = 48 ^ NO 12 - 3 = 9 ^ NO 11 - 4 = 7 ^ NO 11 - 3 = 8 11 * 3 = 33 ^ YES!!!

OpenStudy (anonymous):

3 and 11 because 11-3 = 8 (difference) and 11x3 = 33 (product)